# Concurrence fill and mode distribution of entanglement in neutrino oscillation

### v1

@@ 3 Entanglement Measures in terms of Oscillation Probabilities
The flavor state in Eq. (8) admits several natural bipartitions, allowing one to examine how quantum correlations are distributed among different subsets of flavor modes. To quantify these correlations, the essential tool is the reduced density matrix, obtained by tracing over one or more subsystems of the full state. Starting from the expression of density operator of a specific flavor state,
$|,\rho_{\alpha}(t)=\ket{\nu_{\alpha}(t)}\bra{\nu_{\alpha}(t)}\,$,
one may trace out selected flavor modes to obtain the reduced description relevant for a given partition. In this work, the primary focus is on the single flavor reduced density matrix, obtained by tracing over two of the three flavor modes. This quantity plays a central role because experimentally, neutrino detection is flavor selective, involving interactions of a single flavor at a time, and therefore the single mode density matrix contains all the information required to predict observable event rates. Furthermore, mostly the entanglement measures are constructed from this reduced density matrix provides a direct measure of how strongly the chosen flavor is entangled with the remaining modes. To make this concrete, we consider an initial muon-neutrino state $\nu_{\mu}$, as is relevant for accelerator-based long-baseline experiments such as DUNE. Under three-flavor mixing and propagation, the corresponding time-evolved density matrix takes the explicit form Konwar:2024nrd ; Li:2021epj :
@@ 3.1 Tangle and partial tangle in three-flavor neutrino oscillations
We next intend to introduce different tripartite entanglement observables, such as the tangle and the partial tangle, which quantify the degree of pairwise correlations and their dynamical interplay within the single particle multi mode framework. Within the framework of three flavor neutrino oscillations, with the oscillating neutrinos produced in a definite flavor state at the source, the temporal evolution yields a dynamical redistribution of quantum correlations among the three flavor modes. As a result, one can establish bipartite mode entanglement between the flavor subspaces as a direct consequence of neutrino mixing. This observation naturally motivates us to conduct a detailed, quantitative investigation of sharability of entanglement among different flavor modes during propagation.
These entanglement observables can be directly expressed in terms of the elements of the density matrix given in Eq.(10). This links quantum information theoretic concepts with experimentally observable neutrino oscillation probabilities. The standard three flavor neutrino oscillation probabilities are given in terms of time dependent amplitudes $\widetilde{U}_{\alpha\beta}\big(t,\theta,\delta_{\rm CP}\big)$ as given in Eq.(3) as,
@@
It is already stated in Eq.(8) that the time evolved neutrino states form a single-particle three-mode state corresponding to the propagation of an initial flavor eigenstate $\nu_{\alpha}$ where $\widetilde{U}_{\alpha\beta}$ denotes the time evolved amplitude for the transition $\nu_{\alpha}\to\nu_{\beta}$ which can be expressed in terms of baseline lenghth and energy, including matter effects. A natural way to quantify genuine tripartite entanglement in a three flavor neutrino system is through the tangle, $\tau_{\alpha}$, which measures the residual entanglement associated with a given flavor mode after subtracting all pairwise contributions. Within the three mode framework, the tangle for flavor $\alpha$ is defined as
@@
where $C_{\alpha|\beta\gamma}^{2}$ denotes the bipartitioned concurrence between flavor $\alpha$ and the composite subsystem formed by the remaining two modes, while $C_{\alpha\beta}^{2}$ and $C_{\alpha\gamma}^{2}$ represent the standard pairwise concurrences. This structure captures distribution of the total quantum correlation associated with a particular mode between true tripartite entanglement and bipartite sharing.
@@
where the final equality follows again from probability conservation. Identical arguments show that $\tau_{\mu}=\tau_{\tau}=0$ as well. Thus, the tangle vanishes identically for all flavors throughout the evolution. This result carries an important implication. The three flavor neutrino state cannot host GHZ-type entanglement (for which $\tau_{\alpha}$ is nonzero) but instead lies within the class of separable, biseparable, or W-type states. This motivates the need to examine additional multipartite diagnostics such as partial tangles and concurrence fill, to distinguish among these possibilities and fully characterize how quantum correlations are distributed across the flavor modes.  [^2] The genuine tripartite entanglement $\tau$ and partial tangle for GHZ-type and W-class states are summarized as follows Datta:2018fpp : $\boxed{\begin{aligned} \tau_{\psi}&\neq 0\ \Longrightarrow\ \text{GHZ-type entanglement},\\ \tau_{\psi}&=0,\ \tau_{ij}\neq 0\ \Longrightarrow\ \text{W-type entanglement},\\ \tau_{\psi}&=0,\ \tau_{ij}=0\ \Longrightarrow\ \text{biseparable or separable state.}\end{aligned}}$ (17) .
@@
However, when all three probabilities $P_{\alpha e}$, $P_{\alpha\mu}$, and $P_{\alpha\tau}$ remain nonzero, the system genuinely explores the full three-flavor space, and in this regime the partial tangle no longer coincides with the pairwise concurrence, i.e., $\tau_{\alpha\beta}\neq C_{\alpha\beta}$. This distinction reflects the presence of nontrivial mode sharing among all three flavors and signals the onset of authentic three-mode entanglement redistribution. Thus, within the neutrino oscillation framework, the partial tangles coincide exactly with the corresponding bipartite concurrences, implying that all nontrivial quantum correlations are distributed pairwise among the flavor modes rather than residing in a genuine tripartite component.
@@ 3.2 Concurrence Fill in Terms of Neutrino Oscillation Probabilities
Here, we introduce the concept of concurrence fill $C_{F}$ which offers an intuitive way to visualize and quantify how entanglement is shared among three flavor modes. Unlike the well-known tangle $\tau$, which captures only GHZ-type correlations e.g., $|000\rangle$, $|111\rangle$ states and vanishes for W-type states Dur:2000zz comprising of single excitation subspace using $\rangle, |010|010\rangle$, $|001\rangle$, $C_{F}$ remains non-vanishing for both the cases. Such a measure is more appropriate for both class of GHZ and W-type states that reflects not only the presence of genuine tripartite entanglement but also quantify how that entanglement is distributed among different flavor modes during evolution.
@@
In order to appreciate the concurrence fill, it is probably relevant to introduce a geometrical analogy with the triangle inequality. The geometric triangle inequality states that for any triangle the sum of the lengths of any two sides must be greater than or equal to the length of the remaining side. Analogously, a similar relation has bee established known as the entanglement monogamy inequality. This states that the bipartition entanglement of one subsystem and the rest cannot exceed the sum of the remaining bipartition entanglements in a tripartite system. In the experimental aspect of neutrino oscillation this monogamy relation has already been explored in terms of CKW inequality Banerjee:2024lih . Motivated by this analogy, one can construct a concurrence triangle, where the sides represent the bipartition concurrences among the three partitions of a tripartite quantum system as shown in Fig.1. The perimeter of this concurrence triangle quantifies the total tripartite entanglement shared among the three modes. The area of the concurrence triangle, termed as concurrence fill $C_{F}$, turns out to be a genuine measure of tripartite entanglement.
@@
represents the semiperimeter, analogous to the global entanglement of the system. The extra prefactor $16/3$ appear ensures proper normalization such that $0<C_{F}<1$.
In the context of concurrence triangle as shown in figure 1, when the three sides of the triangle are comparable enclosing a finite area, this finite area signifies genuine tripartite entanglement among different modes. If the length of the shortest side of the concurrence triangle becomes zero, the triangle degenerates into a straight line, corresponding to a completely biseparable state. Furthermore, if any of the three sides vanishes, the system collapses into a fully separable state corresponding to the absence of any entanglement between the different modes. Thus, the concurrence fill $C_{F}$ serves not only as a quantitative measure of the amount of entanglement, but also as an indicator of how that entanglement is distributed among the three flavor modes. As noted earlier, a vanishing three–tangle is characteristic of a $W$-type entangled state. In this case, the concurrence triangle collapses to a line, and the concurrence fill attains the fixed value $C_{F}^{2}=8/9$ for the symmetric $W$ state,
@@
For comparison, the Greenberger-Horne-Zeilinger state,
[…]
exhibits maximal genuine tripartite entanglement, corresponding to the upper limit of the measure with $C_{F}=1$. This contrast highlights the role of $C_{F}$ as a geometric indicator capable of distinguishing between inequivalent classes of three-qubit entanglement. In the context of neutrino oscillations, $C_{F}$ is not constant. Instead, it varies continuously with the oscillation parameters. In such a system, the evolution of $C_{F}(L,E)$ as a function of baseline and energy provides a direct means of tracing the redistribution of flavor coherence and mode entanglement among the three flavor sectors. Using the relation for bipartitioned concurrence square $C_{\alpha|{\beta\gamma}}^{2}(L,E)$ given in Eq.(14), the symmetric factor $Q$ Coffman_2000 is expressed in terms of oscillation probabilities as,
@@
Thus, $Q$ characterizes the overall distribution of bipartitioned concurrences among all flavor modes. Using the above Eq.(24), the concurrnce fill $C^{2}_{F}$ of Eq.(20) can be explicitly written in terms of oscillation probabilities as,
@@ 4 Entanglement Analysis at DUNE using GLoBES
While first oscillation maxima can be accessed by all long baseline neutrino experiments, the second oscillation maxima can be accessed by DUNE (also by T2HK) due to the broad band in beam energy. Furthermore, the second oscillation maxima are particularly important because the interference term dependent on the CP violating phase $\delta$ is enhanced by a factor of $\sim 3$ relative to the first oscillation maxima, offering increased sensitivity to CP violating effects Ghosh:2014rna ; Rout:2020emr . Apart from accessing the second oscillation maxima, the flux factor for DUNE plays an important role in analyzing oscillation probabilities. The DUNE flux, distributed roughly over $0.5-3~\mathrm{GeV}$, therefore provides a unique opportunity to explore quantum-informational aspects of oscillations across different energy ranges accommodating both first and second oscillation maxima. While the neutrino flux in most long-baseline experiments is narrowly peaked around the first oscillation maximum–exhibiting an approximately Gaussian profile shown in green-shaded region in Fig. 2 (right-pannel)–the DUNE beam spans a substantially broader energy window displayed in yellow-shaded region (right-pannel). This wide-band coverage enables a detailed study of flavor evolution not only at the first oscillation maximum but also across several intermediate energy bins and, crucially, at the second oscillation maximum. As a result, DUNE offers an experimentally accessible pathway for probing how bipartite and tripartite entanglement measures evolve across multiple oscillation regimes.
@@ B. Bipartite and Bi-partitioned Concurrence
corresponding to a single excitation in the $\mu$-mode and vacuum in the $e$- and $\tau$-modes. This flavor eigenstate is a coherent superposition of three mass eigenstates, each of which during propagation acquires a dynamical phases $\Delta m_{\mathrm{sol}}^{2}L/(4E)$ or $\Delta m_{\mathrm{atm}}^{2}L/(4E)$. Combined with the PMNS mixing elements, the flavor state at a later time evolves to, as given in Eq.(8), as
@@
The behavior of the entanglement measures like bipartite and bi-partitioned concurrence, defined as in Eq.(14) and Eq.(15) are linked to the energy dependence of the probabilities. The numerical values of the survival and appearance probabilities, $P_{\mu\mu}$, $P_{\mu e}$, and $P_{\mu\tau}$, at selected energies are listed in Table 3, where they explicitly satisfy probability conservation. This is reflected in the left panel of Fig. 2 where we have plotted these probabilities as a function of energies. In Fig. 2, we presented the evolution of oscillation probabilities $P_{\mu e}$ (red curve), $P_{\mu\mu}$ (blue curve), $P_{\mu\tau}$ (yellow curve) as a function neutrino energy which will be relevant for estimation of bipartite and bipartition concurrences. For example, the first oscillation peak for $P_{\mu\tau}$ occurs at $E\simeq 2.6$ GeV while the second oscillation peak occurs at $E\simeq 0.8$ GeV. The vertical dashed lines correspond to specific energies describing maximal mixing of $\nu_{\mu}$ and $\nu_{\tau}$ at $E\simeq 1.7$ GeV, no mixing at $E\simeq 1.3$ GeV while the other two correspond to first oscillation peak $E\simeq 2.6$ GeV and the second oscillation peak at $E\simeq 0.8$ GeV. The crossing of two probability curves $P_{\mu\mu}$ (blue) and $P_{\mu\tau}$ (yellow) at $E\simeq 1.7$ GeV correspond to maximal mixing between the flavors $\nu_{\mu}$ and $\nu_{\tau}$. This will help in understanding the effects of mixing and entanglement for specific energy ranges. The right panel Fig. 2 displays the corresponding appearance probability $P_{\mu e}$, clearly identifying both the first and second oscillation maxima that lie within the accessible DUNE energy window.
@@
For the present analysis, we focus on four representative neutrino energies that capture the essential features of flavor evolution across the DUNE spectrum. Two of these energies, correspond to the first and second oscillation maxima, result in large value of the appearance probability $P_{\mu e}$. The remaining two energies, 1.7 GeV and 1.3 GeV, lie between the two maxima and probe qualitatively different probability configurations. As seen in the left panel of Fig. 2, the point at 1.7 GeV satisfies the condition $P_{\mu\mu}\simeq P_{\mu\tau}$, marking a balanced regime in which disappearance and appearance channels contribute comparably.
@@
In contrast, at 1.3 GeV the probabilities satisfy $P_{\mu e}=P_{\mu\tau}=0$, so that $P_{\mu\mu}=1$, corresponding to a purely $\nu_{\mu}$ configuration with no flavor transition. These probability configurations establish the dynamical backdrop against which the multimode quantum correlations evolve. Collectively, these four energies span the main dynamical regions of flavor evolution, offering natural benchmarks for studying how multimode entanglement varies across the DUNE spectrum. To examine this connection explicitly, we now evaluate the bipartite entanglement measures across the same energy range. The two panels of Fig. 3 provide a detailed picture of how bipartite entanglement evolves in a three-flavor neutrino system as it propagates in vacuum (dashed curves) and in matter (solid curves). The left panel shows the bipartitioned concurrences $C^{2}_{\mu|e\tau}$ (dark blue), $C^{2}_{e|\mu\tau}$ (gray), and $C^{2}_{\tau|\mu e}$ (green), each measuring how strongly one flavor mode is entangled with the remaining two.

### v2

@@ 3 Entanglement Measures in terms of Oscillation Probabilities
The flavor state in Eq. (8) admits several natural bipartitions, allowing one to examine how quantum correlations are distributed among different subsets of flavor modes. To quantify these correlations, the essential tool is the reduced density matrix, obtained by tracing over one or more subsystems of the full state. Starting from the expression of the density operator of a specific flavor state,
$|.\rho_{\alpha}(t)=\ket{\nu_{\alpha}(t)}\bra{\nu_{\alpha}(t)}\,$.
One may trace out selected flavor modes to obtain the reduced description relevant for a given partition. In this work, the primary focus is on the single flavor reduced density matrix, obtained by tracing over two of the three flavor modes. This quantity plays a central role because, experimentally, neutrino detection is flavor-selective, involving interactions with a single flavor at a time; therefore, the single mode density matrix contains all the information required to predict observable event rates. Furthermore, mostly the entanglement measures constructed from this reduced density matrix provide a direct measure of how strongly the chosen flavor is entangled with the remaining modes. To make this concrete, we consider an initial muon-neutrino state, $\nu_{\mu}$, relevant for accelerator-based long-baseline experiments such as DUNE. Under three-flavor mixing and propagation, the corresponding time-evolved density matrix takes the explicit form Konwar:2024nrd ; Li:2021epj :
@@ 3.1 Tangle and partial tangle in three-flavor neutrino oscillations
We next intend to introduce different tripartite entanglement observables, such as the tangle and the partial tangle, which quantify the degree of pairwise correlations and their dynamical interplay within the single-particle multi mode framework. Within the framework of three flavor neutrino oscillations, with the oscillating neutrinos produced in a definite flavor state at the source, the temporal evolution yields a dynamical redistribution of quantum correlations among the three flavor modes. As a result, one can establish bipartite mode entanglement between the flavor subspaces as a direct consequence of neutrino mixing. This observation naturally motivates us to conduct a detailed, quantitative investigation of the entanglement sharability among different flavor modes during propagation.
These entanglement observables can be directly expressed in terms of the elements of the density matrix given in Eq.(10). This links quantum information theoretic concepts to experimentally observable neutrino oscillation probabilities. The standard three flavor neutrino oscillation probabilities are given in terms of time dependent amplitudes $\widetilde{U}_{\alpha\beta}\big(t,\theta,\delta_{\rm CP}\big)$ as given in Eq.(3) as,
@@
It is already stated in Eq.(8) that the time evolved neutrino states form a single-particle three-mode state corresponding to the propagation of an initial flavor eigenstate $\nu_{\alpha}$ where $\widetilde{U}_{\alpha\beta}$ denotes the time evolved amplitude for the transition $\nu_{\alpha}\to\nu_{\beta}$ which can be expressed in terms of baseline length and energy, including matter effects. A natural way to quantify genuine tripartite entanglement in a three flavor neutrino system is through the tangle, $\tau_{\alpha}$, which measures the residual entanglement associated with a given flavor mode after subtracting all pairwise contributions. Within the three mode framework, the tangle for flavor $\alpha$ is defined as
@@
where $C_{\alpha|\beta\gamma}^{2}$ denotes the bipartitioned concurrence between flavor $\alpha$ and the composite subsystem formed by the remaining two modes, while $C_{\alpha\beta}^{2}$ and $C_{\alpha\gamma}^{2}$ represent the standard pairwise concurrences. This structure captures the distribution of the total quantum correlation associated with a particular mode between true tripartite entanglement and bipartite sharing.
@@
where the final equality follows from probability conservation again. Identical arguments show that $\tau_{\mu}=\tau_{\tau}=0$ as well. Thus, the tangle vanishes identically for all flavors throughout the evolution. This result carries an important implication. The three flavor neutrino state cannot host GHZ-type entanglement (for which $\tau_{\alpha}$ is nonzero) but instead lies within the class of separable, biseparable, or W-type states. This motivates the need to examine additional multipartite diagnostics, such as partial tangles and concurrence fill, to distinguish among these possibilities and fully characterize the distribution of quantum correlations across the flavor modes.  [^2] The genuine tripartite entanglement $\tau$ and partial tangle for GHZ-type and W-class states are summarized as follows Datta:2018fpp : $\boxed{\begin{aligned} \tau_{\psi}&\neq 0\ \Longrightarrow\ \text{GHZ-type entanglement},\\ \tau_{\psi}&=0,\ \tau_{ij}\neq 0\ \Longrightarrow\ \text{W-type entanglement},\\ \tau_{\psi}&=0,\ \tau_{ij}=0\ \Longrightarrow\ \text{biseparable or separable state.}\end{aligned}}$ (17) .
@@
However, when all three probabilities $P_{\alpha e}$, $P_{\alpha\mu}$, and $P_{\alpha\tau}$ remain nonzero, the system genuinely explores the full three-flavor space, and in this regime the partial tangle no longer coincides with the pairwise concurrence, i.e., $\tau_{\alpha\beta}\neq C_{\alpha\beta}$. This distinction reflects the presence of nontrivial mode sharing across the three flavors and signals the onset of authentic three-mode entanglement redistribution. Thus, within the neutrino oscillation framework, the partial tangles coincide exactly with the corresponding bipartite concurrences, implying that all nontrivial quantum correlations are distributed pairwise among the flavor modes rather than residing in a genuine tripartite component.
@@ 3.2 Concurrence Fill in Terms of Neutrino Oscillation Probabilities
Here, we introduce the concept of concurrence fill $C_{F}$, which offers an intuitive way to visualize and quantify how entanglement is shared among three flavor modes. Unlike the well-known tangle $\tau$, $C_{F}$ remains non-vanishing for both cases. Such a measure is more appropriate
that reflects not only the presence of genuine tripartite entanglement but also quantifies how that entanglement is distributed among different flavor modes during evolution.
@@
To appreciate the concurrence fill, it is relevant to introduce a geometrical analogy using the triangle inequality. The geometric triangle inequality states that for any triangle, the sum of the lengths of any two sides must be greater than or equal to the length of the remaining side. Analogously, a similar relation, known as the entanglement monogamy inequality, has been established. This states that the bipartition entanglement of one subsystem and the rest cannot exceed the sum of the remaining bipartition entanglements in a tripartite system. In the experimental aspect of neutrino oscillation, this monogamy relation has already been explored in terms of CKW inequality Banerjee:2024lih . Motivated by this analogy, one can construct a concurrence triangle, where the sides represent the bipartition concurrences among the three partitions of a tripartite quantum system as shown in Fig.1. The perimeter of this concurrence triangle quantifies the total tripartite entanglement shared among the three modes. The area of the concurrence triangle, termed as concurrence fill $C_{F}$, turns out to be a genuine measure of tripartite entanglement.
@@
represents the semiperimeter, analogous to the global entanglement of the system. The extra prefactor $16/3$ appears to ensure proper normalization, such that $0<C_{F}<1$.
@@
Thus, $Q$ characterizes the overall distribution of bipartitioned concurrences among all flavor modes. Using the above Eq.(22), the concurrence fill $C^{2}_{F}$ of Eq.(20) can be explicitly written in terms of oscillation probabilities as,
@@ 4 Entanglement Analysis at DUNE using GLoBES
While all long baseline neutrino experiments can access the first oscillation maximum, the second oscillation maximum can be accessed by DUNE (and T2HK) due to the broad band in beam energy. Furthermore, the second oscillation maxima are particularly important because the interference term dependent on the CP violating phase $\delta$ is enhanced by a factor of $\sim 3$ relative to the first oscillation maxima, offering increased sensitivity to CP violating effects Ghosh:2014rna ; Rout:2020emr . Apart from accessing the second oscillation maximum, the flux factor for DUNE plays an important role in analyzing oscillation probabilities. The DUNE flux, distributed roughly over $0.5-3~\mathrm{GeV}$, therefore provides a unique opportunity to explore quantum-informational aspects of oscillations across different energy ranges, accommodating both first and second oscillation maxima. While the neutrino flux in most long-baseline experiments is narrowly peaked around the first oscillation maximum–exhibiting an approximately Gaussian profile shown in green-shaded region in Fig. 2 (right-pannel)–the DUNE beam spans a substantially broader energy window displayed in yellow-shaded region (right-pannel). This wide-band coverage enables a detailed study of flavor evolution not only at the first oscillation maximum but also across several intermediate energy bins and, crucially, at the second oscillation maximum. As a result, DUNE offers an experimentally accessible pathway to probe how bipartite and tripartite entanglement measures evolve across multiple oscillatory regimes.
@@ B. Bipartite and Bi-partitioned Concurrence
corresponding to a single excitation in the $\mu$-mode and vacuum in the $e$- and $\tau$-modes. This flavor eigenstate is a coherent superposition of three mass eigenstates, each of which during propagation acquires a dynamical phase $\Delta m_{\mathrm{sol}}^{2}L/(4E)$ or $\Delta m_{\mathrm{atm}}^{2}L/(4E)$. Combined with the PMNS mixing elements, the flavor state at a later time evolves to, as given in Eq.(8), as
@@
The behavior of the entanglement measures like bipartite and bi-partitioned concurrence, defined as in Eq.(14) and Eq.(15), is linked to the energy dependence of the probabilities. The numerical values of the survival and appearance probabilities, $P_{\mu\mu}$, $P_{\mu e}$, and $P_{\mu\tau}$, at selected energies are listed in Table 3, where they explicitly satisfy probability conservation. This is reflected in the left panel of Fig. 2, where we have plotted these probabilities as a function of energies. In Fig. 2, we presented the evolution of oscillation probabilities $P_{\mu e}$ (red curve), $P_{\mu\mu}$ (blue curve), $P_{\mu\tau}$ (yellow curve) as a function neutrino energy which will be relevant for estimation of bipartite and bipartition concurrences. For example, the first oscillation peak for $P_{\mu\tau}$ occurs at $E\simeq 2.6$ GeV while the second oscillation peak occurs at $E\simeq 0.8$ GeV. The vertical dashed lines correspond to specific energies describing maximal mixing of $\nu_{\mu}$ and $\nu_{\tau}$ at $E\simeq 1.7$ GeV, no mixing at $E\simeq 1.3$ GeV, while the other two correspond to the first oscillation peak at $E\simeq 2.6$ GeV and the second oscillation peak at $E\simeq 0.8$ GeV. The crossing of two probability curves $P_{\mu\mu}$ (blue) and $P_{\mu\tau}$ (yellow) at $E\simeq 1.7$ GeV correspond to maximal mixing between the flavors $\nu_{\mu}$ and $\nu_{\tau}$. This will help in understanding the effects of mixing and entanglement for specific energy ranges. The right panel Fig. 2 displays the corresponding appearance probability $P_{\mu e}$, clearly identifying both the first and second oscillation maxima that lie within the accessible DUNE energy window.
@@
For the present analysis, we focus on four representative neutrino energies that capture the essential features of flavor evolution across the DUNE spectrum. Two of these energies correspond to the first and second oscillation maxima, resulting in a large value of the appearance probability $P_{\mu e}$. The remaining two energies, 1.7 GeV and 1.3 GeV, lie between the two maxima and probe qualitatively different probability configurations. As seen in the left panel of Fig. 2, the point at 1.7 GeV satisfies the condition $P_{\mu\mu}\simeq P_{\mu\tau}$, marking a balanced regime in which disappearance and appearance channels contribute comparably.
@@
In contrast, at 1.3 GeV the probabilities satisfy $P_{\mu e}=P_{\mu\tau}=0$, so that $P_{\mu\mu}=1$, corresponding to a purely $\nu_{\mu}$ configuration with no flavor transition. These probability configurations establish the dynamical backdrop against which the multimode quantum correlations evolve. Collectively, these four energies span the main dynamical regions of flavor evolution, offering natural benchmarks for studying how multimode entanglement varies across the DUNE spectrum. To explicitly examine this connection, we now evaluate bipartite entanglement measures across the same energy range. The two panels of Fig. 3 provide a detailed picture of how bipartite entanglement evolves in a three-flavor neutrino system as it propagates in vacuum (dashed curves) and in matter (solid curves). The left panel shows the bipartitioned concurrences $C^{2}_{\mu|e\tau}$ (dark blue), $C^{2}_{e|\mu\tau}$ (gray), and $C^{2}_{\tau|\mu e}$ (green), each measuring how strongly one flavor mode is entangled with the remaining two.